a) An aqueous solution containing 4% nonvolatile solute exerts a pressure of 1.002 bar at the normal boiling point of the solvent. What is the molar mass of the solute?
b) Why is common salt used to clear the snow off the road? c) Why does water stop boiling when sugar is added to it when it is boiling?

Grade:Upto college level

1 Answers

Sunil Kumar FP
9 years ago
(a)let the molar mass of the solute be x
In 100 gm of the solution ,solute is 4gm and solvent weight that is water is 96 gm mole of the solute =4/x
mole of water=96/18=5.3
molefraction of the solute=(4/x)/5.33+4/x
.lowering of vapor pressure=1.013-1.002 bar=.011 bar
ans relative lowering of vapour pressure=.011/1.013=.011(approx)
this relative lowering is equal to mole fraction of solute=
.011=(4/x)/5.3+4/x
x=67.5gm

(b)This is due to the fact that the melting point of water is decreased on adding salt and water melts and road is being cleared
(c)same above explanation elevation in bioling point